OpenStudy (anonymous):

smallest to biggest? 0.5,1.3,1 1/2,1/5

OpenStudy (anonymous):

0.5, 1.3, \[1\frac{ 1 }{ 2 }=1.5\]\[\frac{ 1 }{ 5 }=0.2\] Thus, the sequence is:\[\frac{ 1 }{ 5 },0.5,1.3,1\frac{ 1 }{ 2 }\]